OpenStudy (anonymous):

Simplify (1 − sin x)(1 + sin x).

OpenStudy (anonymous):

1 cos² x sin² x tan² x

OpenStudy (mathmale):

Hint: (a-b)(a+b) = a^2 - b^2. This is called a "special product."

OpenStudy (mathmale):

Therefore, (1-sin x)(1+sin x) = ?

OpenStudy (anonymous):

(1 - sin x)(1 + cos x) = 1 - sin^2 x = cos^2 x